# Optimal energy storage and collective charging speedup in the central-spin quantum battery

### v1

@@ IV Optimal energy storage and collective charging speedup at non-TC Limits
Having successfully applied the SU(2) symmetry theory to address the optimal energy storage problem in the TC limits, in this section, we will use it again to investigate the problem in the non-TC limits. The non-TC limits distinguish the central-spin QB from the TC QB, representing a unique regime specific to the central-spin QB.
The first non-TC limit considered here is $N_{b}\!\ll\!m\!=\!kN_{c}$ with $k\!\in\!(0,1)$. Since $N_{b}\ll m$, the charger part can be considered essentially unchanged during the dynamics. Therefore, macroscopic excitations are always present in the charger, i.e.,
@@
which prevents the direct application of the HP transformation given in Eq. (III). To circumvent this issue, we can first apply the HP transformation, Eq. (III), and then shift the bosonic operators as follows
@@
Figure 4: (a) The maximum energy storage efficiency $\eta(T)$ as a function of $N_{c}/N_{b}$ for the case $m\!=\!N_{b}$. (b) Charging dynamics for $m\!=\!N_{b}\!=\!N_{c}=100$. The coupling strength is set to $A\!=\!1$.
@@
$=$ $\sqrt{k(1-k)}N_{c}+\mathcal{O}(\sqrt{N_{c}}),$
where we used Eq. (IV). To leading order in $N_{c}$, the central-spin model in Eq. (II) reduces to
$H$ $\overset{N_{b}\!\ll\!m=kN_{c}}{\simeq}$ $A\sqrt{k(1-k)}N_{c}(S^{-}+S^{+})$ (29)
@@
which is exactly the generator of the SU(2) algebra, and therefore means that SU(2) symmetry also appears in this non-TC limit. The evolution of energy storage (4) in this case can be expressed as
$\Delta E(t)=\frac{N_{b}\omega}{2}+\langle\psi(0)|S^{z}(t)|\psi(0)\rangle,$ (30)
@@
Note that Eq. (33) does not hold for the case $k\!=\!1$, or equivalently $m\!=\!N_{c}$, because the approximation of $J^{+}$ is incorrect in this scenario, as the zeroth-order term vanishes. To investigate the region where $m\!=\!N_{c}$, we perform numerical calculations, see Fig. (4). In Fig. 4(a) we obtain the maximum energy transferred into the battery and plot the energy storage efficiency $\eta(T)$, defined in Eq. (5), at the charging time $T$ for various values of $N_{b}$. Surprisingly, $\eta(T)$ exhibits a universal behavior, with all curves collapsing onto a single curve after rescaling. Moreover, we observe that $\eta(T)\!\to\!1$ for the case $N_{c}\!=\!N_{b}$, indicating the possibility of achieving optimal energy storage. In Fig. 4(b), we plot the charging dynamics for $N_{b}\!=\!m\!=\!N_{c}$ which shows asymptotic optimal energy storage, meaning that the first maximum of $\eta$ approaches one. However, due to the absence of dynamically emergent SU(2) symmetry, this case differs significantly from the previous situations and does not exhibit sine-like oscillations. In Fig. 5, we plot the collective charging advantage for $N_{b}\!=\!m\!=\!N_{c}$ and find that $\Gamma\propto N_{b}^{0.8264}$. Although $\Gamma$ does not scale exactly with $N_{b}$ in this case, it still demonstrates a collective charging advantage close to an $N_{b}$-scaling behavior.
@@
Cases Charging Time Optimal Energy Storage Emergence SU(2) Symmetry Collective Charging Advantage
@@
V Conclusion
@@
In this work, we investigate the charging dynamics of a central-spin QB. The problems of optimal energy storage and collective charging advantage are analytically explored within this model. We rigorously demonstrate the realization of optimal energy storage in four different initial state settings: 1) $N_{b}\!\ll\!m\!\ll\!N_{c}$; 2) $m\!\ll\!N_{b}\!\ll\!N_{c}$; 3) $m\!\ll\!N_{c}\!\ll\!N_{b}$; 4) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$], where $N_{b}$ denotes the number of battery cells, $N_{c}$ represents the number of charger units, and $m$ is the initial number of excited charger units. The first three cases correspond to the TC limits, while the fourth case represents the non-TC limit. Our findings reveal that, across all these cases, the emergence of SU(2) symmetry from the charging dynamics is crucial for achieving optimal energy storage. This emergent symmetry also provides a theoretical approach to analytically determine the charging time. In the fully charged scenarios corresponding to cases (1) and (4), a collective charging speedup of $N_{b}$-fold is proven in comparison to the parallel charging scheme. Furthermore, in the case of $N_{b}\!=\!m\!=\!N_{c}$ we identify the possibility of asymptotically achieving optimal energy storage, with numerical calculations indicating a collective charging advantage scaling as approximately $\sim N_{b}^{0.8264}$. The main results are summarized in Table 1.
@@
Our work highlights the significance of emergent SU(2) symmetry in analytically addressing the challenges of optimal energy storage and collective charging speedup. It extends previous studies on the TC QB Yang et al. (2024) and the central-spin QB Liu et al. (2021). Especially in cases (1) and (4), not only the optimal energy storage can be achieved, but also the collective charging advantage of $N_{b}$ exists, which provides guidance for designing high-performance QBs. We believe the methods used here can be applied to other QB models to uncover further the mechanisms underlying optimal energy storage and collective charging speedup.
@@ ACKNOWLEDGMENTS
This work was supported by the NSFC (Grants No. 12275215, No. 12305028, and No. 12247103), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSZ005) and the Youth Innovation Team of Shaanxi Universities. HLS was supported by the European Commission through the H2020 QuantERA ERA-NET Cofund in Quantum Technologies project “MENTA”.

### v2

@@ IV Optimal energy storage and collective charging speedup at non-TC Limits
Having successfully applied the SU(2) symmetry theory to address the optimal energy storage problem in the TC limits, in this section we use it again to investigate the problem in the non-TC limits. The non-TC limits distinguish the central-spin QB from the TC QB, representing a unique regime specific to the central-spin QB.
The first non-TC limit considered here is $N_{b}\!\ll\!m\!=\!kN_{c}$, with $k\!\in\!(0,1)$. Since $N_{b}\ll m$, the charger part can be considered essentially unchanged during the dynamics. Therefore, macroscopic excitations are always present in the charger, i.e.,
@@
which prevents the direct application of the HP transformation given in Eq. (III). To circumvent this issue, we can first apply the HP transformation (III) and then shift the bosonic operators as
@@
Figure 4: (a) Maximum energy storage efficiency $\eta(T)$ as a function of $N_{c}/N_{b}$ for the case $m\!=\!N_{b}$. (b) Charging dynamics for $m\!=\!N_{b}\!=\!N_{c}=100$. The coupling strength is set to $A\!=\!1$.
@@
$=$ $\sqrt{k(1-k)}N_{c}+O(\sqrt{N_{c}}),$
where we have used Eq. (IV). To leading order in $N_{c}$, the central-spin model in Eq. (II) reduces to
$H$ $\overset{N_{b}\!\ll\!m=kN_{c}}{\longrightarrow}$ $A\sqrt{k(1-k)}N_{c}(S^{-}+S^{+})$ (32)
@@
which is exactly the generator of the SU(2) algebra and therefore means that SU(2) symmetry also appears in this non-TC limit. The evolution of energy storage (7) in this case can be expressed as
$\Delta E(t)=\frac{N_{b}\omega}{2}+\langle\psi(0)|S^{z}(t)|\psi(0)\rangle\omega,$ (33)
@@
Note that Eq. (36) does not hold for the case $k\!=\!1$, or equivalently $m\!=\!N_{c}$, because the approximation of $J^{+}$ is incorrect in this scenario, as the zeroth-order term vanishes. To investigate the region where $m\!=\!N_{c}$, we perform numerical calculations (see Fig. 4). In Fig. 4(a) we obtain the maximum energy transferred into the battery and plot the energy storage efficiency $\eta(T)$, defined in Eq. (8), at the charging time $T$ for various values of $N_{b}$. Surprisingly, $\eta(T)$ exhibits a universal behavior, with all curves collapsing onto a single curve after rescaling. Moreover, we observe that $\eta(T)\!\to\!1$ for the case $N_{c}\!=\!N_{b}$, indicating the possibility of achieving optimal energy storage. In Fig. 4(b) we plot the charging dynamics for $N_{b}\!=\!m\!=\!N_{c}$, which shows asymptotic optimal energy storage, meaning that the first maximum of $\eta$ approaches one. However, due to the absence of dynamically emergent SU(2) symmetry, this case differs significantly from the previous situations and does not exhibit sine like oscillations. In Fig. 5 we plot the collective charging advantage for $N_{b}\!=\!m\!=\!N_{c}$ and find that $\Gamma\propto N_{b}^{0.8264}$. Although $\Gamma$ does not scale exactly with $N_{b}$ in this case, it still demonstrates a collective charging advantage close to an $N_{b}$-scaling behavior.
V Origin of collective charging advantage
We emphasize that there are two distinct origins of the collective charging advantage: the classical origin [36, 12, 7] and the quantum origin [11]. The classical origin arises from the scaling amplification of the norm of the collective charging Hamiltonian, rather than from quantum correlations, such as multipartite entanglement. We argue that the collective charging advantage observed in the central-spin quantum battery is also classical in nature, similar to that of the Dicke battery [13].
The first piece of evidence comes from the quantum speed limit [37, 38]
$T\geq\frac{\pi\hbar}{2\Delta E},$ (38)
where $T$ is the charging time, $\hbar\!=\!1$ in all calculations, and $\Delta E\!=\!\sqrt{\langle\psi(0)|H^{2}|\psi(0)\rangle-\langle\psi(0)|H|\psi(0)\rangle^{2}}$. If the charging time saturates the quantum speed limit, then the charging time $T^{\prime}$ generated by the dynamics $H^{\prime}\!=\!N_{b}H$ will be $1/N_{b}$ times the original charging time $T$, which is a purely classical effect. Therefore, we introduce the charging rate defined as
$\gamma\equiv\frac{\tau_{\rm QSL}}{T}$ (39)
to witness the classical origin of the collective charging advantage, where $\tau_{\rm QSL}=\pi/2\Delta E$.
A typical parallel charging scheme is governed by Hamiltonian $H_{\parallel}\!=\!S^{x}\!=\!\sum_{j=1}^{N_{b}}\sigma_{j}^{x}/2$. It follows that $T_{\parallel}\!=\!\pi$ and $\Delta E_{\parallel}=\sqrt{N_{b}}/2$. Thus, the typical charging rate for a parallel charging scheme is
$\gamma_{\parallel}=\frac{1}{\sqrt{N_{b}}}.$ (40)
For the collective charging scheme in the central-spin QB, the variance of charging Hamiltonian (II) is given by
$\Delta E$ $=$ $A\sqrt{\langle 0|S^{-}S^{+}|0\rangle\langle m|J^{+}J^{-}|m\rangle}$ (41)
$=$ $A\sqrt{N_{b}m(N_{c}-m+1)}.$
Now we consider the charging rate for two cases with collective charging advantage. The first case corresponds to the TC limit discussed in Sec. III, i.e., $N_{b}\!\ll\!m\!\ll\!N_{c}$. Substituting the charing time (16) into the definition of charging rate (39), we obtain
$\gamma(N_{b}\!\ll\!m\!\ll\!N_{c})\!=\!\frac{\sqrt{N_{c}}\sqrt{m-\frac{N_{b}-1}{2}}}{\sqrt{N_{b}m(N_{c}-m+1)}}\simeq\frac{1}{\sqrt{N_{b}}}.$ (42)
The second case corresponds to the non-TC limit $N_{b}\!\ll\!m\!=\!kN_{c}$ discussed in Sec. IV. By substituting the charing time (36) into the charging rate (39), we have
$\gamma(N_{b}\!\ll\!m\!=\!kN_{c})\!=\!\frac{N_{c}\sqrt{k(1-k)}}{\sqrt{N_{b}m(N_{c}-m+1)}}\!\simeq\!\frac{1}{\sqrt{N_{b}}}.$ (43)
Since the collective charging rate in the central-spin quantum battery, given by Eqs. (42) and (43), is the same as the parallel charging rate in Eq. (40), we conclude that the collective charging advantage in the central-spin quantum battery is also classical due to the super linear scaling of the bandwidth of the charging Hamiltonian.
Figure 6: Evolution of the rescaled QFI for the battery state in the cases of $N_{b}\!\ll\!m\!\ll\!N_{c}$ and $N_{b}\!\ll\!m\!=\!kN_{c}$, in which an $N_{b}$-fold enhancement in speedup exists. The number of battery cells is set to $N_{b}\!=\!50$.
@@ V Origin of collective charging advantage
Case Charging Time Optimal Energy Storage Emergence SU(2) Symmetry Collective Charging Advantage
@@
The second piece of evidence comes from analyzing the multipartite entanglement structure during the charging dynamics. To witness multipartite entanglement, we adopt the quantum Fisher information (QFI). For a generic $N$-qubit state $\rho\!=\!\sum_{\lambda}q_{\lambda}|\psi_{\lambda}\rangle\langle\psi_{\lambda}|$ (in diagonal form) and an operator $\hat{\mathcal{O}}$, the QFI is given by [39]
$F_{Q}[\rho,\hat{\mathcal{O}}]\!=\!2\sum_{\lambda,\lambda^{\prime}}\frac{(q_{\lambda}-q_{\lambda^{\prime}})^{2}}{q_{\lambda}+q_{\lambda^{\prime}}}|\langle\psi_{\lambda}|\hat{\mathcal{O}}|\psi_{\lambda^{\prime}}\rangle|^{2}.$ (44)
Violation of the inequality
$F_{Q}[\rho_{b}(t),S^{x}]\!\leq\!sk^{2}+r^{2},$ (45)
for $\hat{\mathcal{O}}\!=\!S^{x}$ indicates that the battery state $\rho_{b}(t)$ contains at least $(k\!+\!1)$-particle entanglement among the $N_{b}$ battery cells [40, 41, 42], where $s\!=\!\lfloor{N_{b}/k}\rfloor$ represents the largest integer less than or equal to $N_{b}/k$, and $r\!=\!N_{b}\!-\!sk$. Therefore, if $F_{Q}[\rho_{b}(t),S^{x}]\!>\!N_{b}$, we can infer the appearance of an entangled battery state. Conversely, if $F_{Q}[\rho_{b}(t),S^{x}]\!\leq\!N_{b}$, we cannot extract useful information about the entanglement structure of $\rho_{b}(t)$. Actually, QFI is particularly useful for detecting multipartite entangled states close to the Greenberger–Horne–Zeilinger (GHZ)-like states, as the QFI for the GHZ state is $N_{b}^{2}$, larger than $N_{b}$.
Recall the results obtained in the previous sections: We have shown that the appearance of optimal charging is associated with the emergent SU(2) symmetry. In particular, for cases with collective charging advantage, the Hamiltonian in Eq. (2) can be approximated as a generator of SU(2) algebra, i.e.,
$H\simeq\nu S^{x},$ (46)
where $\nu=2A\sqrt{N_{c}}\sqrt{m-(N_{b}-1)/2}$ for the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$ [43], while $\nu=2A\sqrt{k(1-k)}N_{c}$ for the case of $N_{b}\!\ll\!m\!=\!kN_{c}$, as seen in Eq. (32). Based on the effective Hamiltonian in Eq. (46), the evolution of the component of the wave function (4) is obtained as
$\bm{\psi}_{m}(t)\!=\!\sqrt{\binom{N_{b}}{m}}\left[-i\sin\left(\frac{\nu}{2}t\right)\right]^{m}\left[\cos\left(\frac{\nu}{2}t\right)\right]^{N_{b}-m}.$ (47)
Therefore, using Eq. (6), the reduced state of the battery is given by
$\rho_{b}(t)$ $=$ $\sum_{m=0}^{N_{b}}q_{m}(t)|m\rangle\langle m|,$ (48)
with eigenvalues
[…]
Substituting Eq. (48) into the definition of QFI (44), we obtain
$F_{Q}[\rho_{b}(t),S^{x}]$
$=\sum_{m=0}^{N_{b}-1}\frac{[q_{m}(t)-q_{m+1}(t)]^{2}}{q_{m}(t)+q_{m+1}(t)}(N_{b}-m)(m+1),$ (50)
where $q_{m}(t)$ is given in Eq. (49). Figure 6 plots the evolution of QFI for the battery state, given by Eq. (V), which is always less than or equal to $N_{b}$. Therefore, a GHZ-like multipartite entanglement structure has not been detected among the battery cells using QFI. With these two pieces of evidence, we conclude that the collective charging advantage observed in the central-spin quantum battery has no genuine quantum origin.
VI Conclusion
@@
In this work we investigated the charging dynamics of a central-spin QB. The problems of optimal energy storage and collective charging advantage were analytically explored within this model. We rigorously demonstrated the realization of optimal energy storage in four different initial state settings: (i) $N_{b}\!\ll\!m\!\ll\!N_{c}$, (ii) $m\!\ll\!N_{b}\!\ll\!N_{c}$, (iii) $m\!\ll\!N_{c}\!\ll\!N_{b}$, and (iv) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$], where $N_{b}$ denotes the number of battery cells, $N_{c}$ represents the number of charger units, and $m$ is the initial number of excited charger units. The first three cases correspond to the TC limits, while the fourth case represents the non-TC limit. Our findings reveal that, across all these cases, the emergence of SU(2) symmetry from the charging dynamics is crucial for achieving optimal energy storage. This emergent symmetry also provides a theoretical approach to analytically determine the charging time. In the fully charged scenarios corresponding to cases (i) and (iv), a collective charging speedup of $N_{b}$-fold is proven in comparison to the parallel charging scheme. Furthermore, in the case of $N_{b}\!=\!m\!=\!N_{c}$ we identified the possibility of asymptotically achieving optimal energy storage, with numerical calculations indicating a collective charging advantage scaling as approximately $\sim N_{b}^{0.8264}$. The main results are summarized in Table 1. By employing the quantum speed limit and multipartite entanglement witness, we identified the classical origin of the collective charging advantage in the central-spin quantum battery. Nevertheless, the classical collective charging advantage is also valuable for designing high-performance quantum batteries [7, 44]. The theoretical central-spin quantum battery in Eq. (II) can be realized in a strongly interacting dipolar spin ensemble [45], where a high-spin NV center can effectively correspond to battery cells, and an ensemble of spin-1/2 P1 centers represent charging units. The key flip-flop spin interactions $H_{I}$ in Eq. (II) can be realized through the magnetic dipole-dipole interaction between NV and P1 centers. From a practical viewpoint, the case where $N_{b}\!=\!m\!=\!N_{c}$ is the most experimentally favorable for demonstrating both optimal energy storage and the collective charging advantage.
@@
Our work highlights the significance of emergent SU(2) symmetry in analytically addressing the challenges of optimal energy storage and collective charging speedup. It extends previous studies on the TC QB [16] and the central-spin QB [35]. We believe the methods used here can be applied to other QB models to uncover further the mechanisms underlying optimal energy storage and collective charging speedup.
@@ ACKNOWLEDGMENTS
HLS thanks Xu Zhou for providing valuable experimental comments. This work was supported by the NSFC (Grants No. 12275215, No. 12305028, and No. 12247103), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSZ005) and the Youth Innovation Team of Shaanxi Universities. HLS was supported by the European Commission through the H2020 QuantERA ERA-NET Cofund in Quantum Technologies project “MENTA”.